# Faster Linear Systems and Matrix Norm Approximation via Multi-level Sketched Preconditioning

### v1

@@ Abstract
We present a new class of preconditioned iterative methods for solving linear systems of the form $\mathbf{A}\mathbf{x}=\mathbf{b}$. Our methods are based on constructing a low-rank Nyström approximation to $\mathbf{A}$ using sparse random sketching. This approximation is used to construct a preconditioner, which itself is inverted quickly using additional levels of random sketching and preconditioning.
@@
    We give the first $\tilde{O}(n^{2}+{d_{\lambda}}^{\omega}$) time algorithm for solving a regularized linear system $(\mathbf{A}+\lambda\mathbf{I})\mathbf{x}=\mathbf{b}$, where $\mathbf{A}$ is positive semidefinite with effective dimension $d_{\lambda}$. This problem arises in applications like Gaussian process regression.
@@
    We give faster algorithms for approximating Schatten $p$-norms and other matrix norms. For example, for the Schatten 1 (nuclear) norm, we give an algorithm that runs in $\tilde{O}(n^{2.11})$ time, improving on an $\tilde{O}(n^{2.18})$ method of [Musco et al., ITCS 2018].
Interestingly, previous state-of-the-art algorithms for most of the problems above relied on stochastic iterative methods, like stochastic coordinate and gradient descent. Our work takes a completely different approach, instead leveraging tools from matrix sketching.
@@ 1 Introduction
We consider the complexity of solving a system of linear equations: given an $n\times n$ matrix $\mathbf{A}$ and an $n$-dimensional vector $\mathbf{b}$, find $\mathbf{x}$ such that $\mathbf{A}\mathbf{x}=\mathbf{b}$. This ubiquitous task in numerical linear algebra has applications across many areas, from data science and machine learning, to physical sciences, engineering, and more. Despite many existing methods for solving linear systems, they remain a major computational bottleneck, so faster algorithms are a subject of active research. Much of the progress on faster algorithms concentrates on designing methods tailored for matrices with special structure, such as Laplacian, Toeplitz, or Hankel matrices, among others [KKM79, XXG12, ST14, KMP12, KS16, CKK+18, PV21]. Progress on general, unstructured linear systems has been slower.
@@
When $k=n$, our fastest methods for solving 1.1 run in $O(n^{\omega})$ time via fast matrix multiplication. When $k=0$, the problem can be solved in $\tilde{O}(n^{2})$ time via any standard iterative solver such as CG. We are interested in the arguably more interesting intermediate regime, where $\mathbf{A}$ is a combination of a low-rank ill-conditioned matrix and a full-rank well-conditioned one. Such matrices arise often in practice, either due to various forms of regularization that are prevalent in machine learning, statistics and optimization [BV04, ZDW13, DW18], or due to the presence of isotropic noise coming from measurement error, rounding, or compression [LW11, LGW+21].
A standard approach to Problem 1.1 is to construct an approximation to the rank $k$ subspace identified by $\mathbf{A}$’s top singular values. That information can be used to precondition an iterative solver so that it runs in $\tilde{O}(n^{2})$ time [HMT11]. Constructing a sufficiently accurate approximation to the top subspace requires (block) power iteration or related methods [MM15], which take $\tilde{O}(n^{\omega(\log_{n}k)})$ time, where $\omega(\log_{n}k)\in[2,\omega]$ is the exponent of rectangular matrix multiplication between an $n\times n$ and $n\times k$ matrix [LG12]. [^1] This corresponds to the $O(n^{2}k)$ time for classical multiplication of $n\times n$ and $n\times k$ matrices. This approach already interpolates between the two extremes of the problem. In fact, it achieves a near-optimal $\tilde{O}(n^{2})$ runtime for $k=O(n^{0.32})$.
@@ 1.1 Main Results
$\tilde{O}\left(\left(n^{2.065}+k^{\omega}\right)\cdot\log^{3}1/\epsilon\right).$
@@
Surprisingly, our algorithm departs from the stochastic optimization approach used in the prior work of [DY24]. Instead, we employ a deterministic solver with a randomized preconditioner. We build on two key observations. The first, which has been made in prior work [FTU23, DEF+23], is that even a coarse approximation to the top rank $k$ subspace can be used to construct a “good enough” preconditioner for $\mathbf{A}$. This fact allows us to replace the use of algorithms like block power method with more efficient linear-time sketching methods for low-rank approximation [CW13]. Second, and more importantly, we show that it is possible to avoid explicitly constructing a rank $k$ subspace approximating the top singular vectors of $\mathbf{A}$, which would cost $\tilde{O}(nk^{\omega-1})$ time. Instead, using $\tilde{O}(nk+k^{\omega})$ preprocessing time, we maintain a data structure that lets us perform inexact matrix-vector products with an inverted low-rank preconditioner. The data structure involves additional levels of solving linear systems with randomized sketching and preconditioning methods, resulting in a framework that we call Multi-level Sketched Preconditioning (MSP).
We discuss details of our approach further in Section 1.2. Before doing so, we highlight two additional applications of the framework beyond Theorem 1.1.
@@
An important special case of linear systems that are well-conditioned except for a few large singular values are those that are explicitly regularized by adding a scaled identity, $\lambda\mathbf{I}$, to a positive semidefinite matrix $\mathbf{A}$, a setting which arises, for instance, in kernel ridge regression [EAM14, RCR17, MM17, ACW17, FTU23] and second-order optimization [Lev44, Mar63, MS12, JJM23]. Prior results have shown that such linear systems can be solved in $\tilde{O}(n^{2}+n{d_{\lambda}}^{\omega-1})$, where $d_{\lambda}={\operatorname{tr}}(\mathbf{A}(\mathbf{A}+\lambda\mathbf{I})^{-1})\leq n$ is the so-called $\lambda$-effective dimension of the problem, often much smaller than $n$ [ACW17, MM17]. Our algorithms improve on this by using Multi-level Sketched Preconditioning, leading to a time complexity of $\tilde{O}(n^{2}+{d_{\lambda}}^{\omega})$, which is optimal (up to log factors).
@@
$\tilde{O}\left(\left(n^{2}+{d_{\lambda}}^{\omega}\right)\cdot\log^{3}1/\epsilon\right),\quad\text{where}\quad d_{\lambda}={\operatorname{tr}}(\mathbf{A}(\mathbf{A}+\lambda\mathbf{I})^{-1}).$
@@
Linear systems often arise in algorithms for solving other matrix problems, including linear and semidefinite programming [CLS21, JKL+20], least squares and $l_{p}$ regression [Sar06, RT08, MM13, CP15, CD21], as well as for estimating various matrix properties [MNS+18]. Our new algorithms can be used to speed up any of these tasks. As a motivating example, we show how to improve methods for estimating various matrix norms, a fundamental problem in linear algebra. We build on an approach of [MNS+18], which uses regularized linear system solves to build rational function approximations that, when combined with stochastic trace estimation methods [Hut90], allow for the estimation of various functions of the singular values faster than $O(n^{\omega})$ time, i.e. faster than the time it takes to compute all singular values outright. [^3] Concretely, [MNS+18] estimates quantities of the form ${\operatorname{tr}}(f(\mathbf{A}))$, with the nuclear norm equal to the case where $f(x)=|x|$. Stochastic trace estimation methods like Hutchinson’s estimator and related techniques approximate ${\operatorname{tr}}(f(\mathbf{A}))$ by computing $\mathbf{g}^{T}f(\mathbf{A})\mathbf{g}$ for randomly chosen $\mathbf{g}$ [MMMW21].
@@ 1.2 Our Techniques
The starting point for our Multi-level Sketched Preconditioning comes from work on solving regularized systems of the form $(\mathbf{A}+\lambda\mathbf{I})\mathbf{x}=\mathbf{b}$, where $\mathbf{A}$ is positive semidefinite (PSD). Here, a central technique for preconditioning is the Nyström method [Nys30, WS01, GM16, MM17], which uses randomized sketching or sub-sampling to build a low-rank approximation for $\mathbf{A}$. As an illustration, let $\mathbf{S}\in\mathbb{R}^{l\times n}$ be a subsampling matrix, so that $\mathbf{A}\mathbf{S}^{\top}$ contains a random subset of $l$ columns from $\mathbf{A}$. The classical Nyström approximation is defined as $\hat{\mathbf{A}}_{\mathrm{nys}}=\mathbf{C}\mathbf{W}^{-1}\mathbf{C}^{\top}$, where $\mathbf{C}=\mathbf{A}\mathbf{S}^{\top}$ is the column submatrix and $\mathbf{W}=\mathbf{S}\mathbf{A}\mathbf{S}^{\top}$ is the corresponding $l\times l$ principal submatrix of $\mathbf{A}$. This can be extended to other sketching methods by replacing $\mathbf{S}$ with a sparse random matrix (CountSketch, OSNAP, etc. [CW13, NN13, Coh16]) or a Subsampled Randomized Hadamard Transform (SRHT, [AC09, Tro11]), all of which can be multiplied by $\mathbf{A}$ in $\tilde{O}(n^{2})$ time. With sufficiently large sketch size (namely, proportional to the effective dimension $d_{\lambda}={\operatorname{tr}}(\mathbf{A}(\mathbf{A}+\lambda\mathbf{I})^{-1})$), one can show that $\mathbf{M}=\hat{\mathbf{A}}_{\mathrm{nys}}+\lambda\mathbf{I}$ is a good preconditioner for $\mathbf{A}+\lambda\mathbf{I}$, i.e., the condition number of $\mathbf{M}^{-1}\mathbf{A}$ is constant. At the same time, $\mathbf{M}^{-1}$ can be applied quickly due to its decomposition into $\mathbf{C}$ and $\mathbf{W}$. Concretely, using the inversion formula $\mathbf{M}^{-1}=\frac{1}{\lambda}\big{(}\mathbf{I}-\mathbf{C}(\mathbf{C}^{\top}\mathbf{C}+\lambda\mathbf{W})^{-1}\mathbf{C}^{\top}\big{)}$, we can compute $\mathbf{M}^{-1}\mathbf{r}$ for any vector $\mathbf{r}$ in $\tilde{O}(nl)$ time after precomputing $(\mathbf{C}^{\top}\mathbf{C}+\lambda\mathbf{W})^{-1}$ in $\tilde{O}(nl^{\omega-1})$ time.
@@
Average Condition Number Bounds via Nyström Preconditioning. We improve on standard Nyström preconditioning in two important ways. First, we show that, even when solving an unregularized system $\mathbf{A}\mathbf{x}=\mathbf{b}$, a preconditioner of the form $\mathbf{M}=\hat{\mathbf{A}}_{\mathrm{nys}}+\lambda\mathbf{I}$, where $\lambda$ is now just a parameter of the algorithm, can significantly improve conditioning. Specifically, we prove that, using a Nyström approximation with rank $\tilde{O}(l)$ for a matrix $\mathbf{A}$ with singular values $\sigma_{1}\geq\ldots\geq\sigma_{n}$, the condition number of the preconditioned system $\mathbf{M}^{-1}\mathbf{A}$ can be reduced to $\tilde{O}(\frac{n}{l}\cdot\bar{\kappa}_{l})$, where $\bar{\kappa}_{l}=\frac{1}{n-l}\sum_{i>l}\frac{\sigma_{i}}{\sigma_{n}}$ is the average condition number of $\mathbf{A}$, excluding its top $l$ singular values. If $\mathbf{A}$ has at most $k$ large singular values, as in 1.1, then $\bar{\kappa}_{l}=O(1)$ for any $l\geq k$. Due to the dependence on $\frac{n}{\ell}$, increasing $l$ past $k$ still improves the preconditioner, but at the expense of increasing the $\tilde{O}(nl^{\omega-1})$ cost of inverting the Nyström approximation. This limitation motivates our second contribution.
@@
Two-level Preconditioning for PSD Systems. Instead of computing $\mathbf{M}^{-1}=\frac{1}{\lambda}\big{(}\mathbf{I}-\mathbf{C}(\mathbf{C}^{\top}\mathbf{C}+\lambda\mathbf{W})^{-1}\mathbf{C}^{\top}\big{)}$ explicitly, we focus on simply implementing matrix-vector multiplications with $\mathbf{M}^{-1}$, which suffice to to apply a preconditioned iterative method like CG or Lanczos. The main challenge in doing so is to implement efficient multiplications with $(\mathbf{C}^{\top}\mathbf{C}+\lambda\mathbf{W})^{-1}$, i.e. to solve the system $(\mathbf{C}^{\top}\mathbf{C}+\lambda\mathbf{W})\mathbf{x}=\mathbf{r}$. We show how to do this using a second level of sketching and preconditioning. In particular, since $\mathbf{C}$ is an $n\times l$ matrix, standard results (e.g., [RT08, CW13]) show that $(\mathbf{C}^{\top}\mathbf{C}+\lambda\mathbf{W})\mathbf{x}=\mathbf{r}$ can be solved efficiently by preconditioning with $\mathbf{C}^{\top}{\mathbf{\Phi}}^{\top}{\mathbf{\Phi}}\mathbf{C}+\lambda\mathbf{W}$, where $\mathbf{\Phi}$ is an $O(l)\times n$ sketching matrix. $\mathbf{C}^{\top}{\mathbf{\Phi}}^{\top}{\mathbf{\Phi}}\mathbf{C}+\lambda\mathbf{W}$ can be constructed and inverted in just $\tilde{O}(nl+l^{\omega})$ time, which leads to an upfront cost of just $\tilde{O}(nl+l^{\omega})$ for implementing $\tilde{O}(nl)$ time matrix-vector multiplications with $\mathbf{M}^{-1}$. This is a significant improvement on the $\tilde{O}(nl^{\omega-1})$ required by explicit inversion. Overall, we obtain a complexity of $\tilde{O}(n^{2}\sqrt{n/l}+l^{\omega})$ for solving 1.1 when $\mathbf{A}$ is PSD. Optimizing over $l$ gives our final result for solving 1.1 in $\tilde{O}(n^{2.065}+k^{\omega})$ time for positive definite $\mathbf{A}$. See Section 4, Theorem 4.1, for details.
We note that implementing the approach above requires a careful analysis of the preconditioned Lanczos method to ensure that an optimal convergence rate can still be obtained even when the preconditioner $\mathbf{M}^{-1}$ is applied inexactly. We give this analysis in Section 6, building on prior work on understanding the stability of the unpreconditioned Lanczos method, and the related CG method, for solving linear systems [Pai71, Pai76, Gre89, MMS18].
@@
Three-level Preconditioning for non-PSD Systems. The above strategy, which involves two levels of preconditioning, is restricted to PSD matrices. Extending the approach to arbitrary linear systems requires additional work. A natural first attempt is to reduce an arbitrary $\mathbf{A}\mathbf{x}=\mathbf{b}$ linear system to a PSD linear system via the normal equations, $\mathbf{A}^{\top}\mathbf{A}\mathbf{x}=\mathbf{A}^{\top}\mathbf{b}$, where $\mathbf{A}^{\top}\mathbf{A}$ is PSD. However, if we apply the Nyström preconditioning approach discussed above, we realize that we require matrices $\mathbf{C}=\mathbf{A}^{\top}\mathbf{A}\mathbf{S}^{\top}$ and $\mathbf{W}=\mathbf{S}\mathbf{A}^{\top}\mathbf{A}\mathbf{S}^{\top}$, which can no longer be computed in $\tilde{O}(n^{2})$ time, even if $\mathbf{S}$ is a sparse random matrix or SRHT sketch. To address this, we show that the matrix $(\mathbf{C}^{\top}\mathbf{C}+\lambda\mathbf{W})^{-1}$ can nevertheless be applied efficiently by preconditioning with the approximation $(\mathbf{W}^{2}+\lambda\mathbf{W})^{-1}=\frac{1}{\lambda}\big{(}\mathbf{W}^{-1}-(\mathbf{W}+\lambda\mathbf{I})^{-1}\big{)}$. This preconditioner, in turn, can be applied by solving two linear systems associated with $\mathbf{W}^{-1}$ and $(\mathbf{W}+\lambda\mathbf{I})^{-1}$, both using a third level of sketching and preconditioning. Details are included in Section 5.

### v2

@@ Abstract
We present a new class of preconditioned iterative methods for solving linear systems of the form $\mathbf{A}\mathbf{x}=\mathbf{b}$. Our methods are based on constructing a low-rank Nyström approximation to $\mathbf{A}$ using sparse random matrix sketching. This approximation is used to construct a preconditioner, which itself is inverted quickly using additional levels of random sketching and preconditioning.
@@
    We give the first $\tilde{O}(n^{2}+{d_{\lambda}}^{\omega}$) time algorithm for solving a regularized linear system $(\mathbf{A}+\lambda\mathbf{I})\mathbf{x}=\mathbf{b}$, where $\mathbf{A}$ is positive semidefinite with effective dimension $d_{\lambda}={\operatorname{tr}}(\mathbf{A}(\mathbf{A}+\lambda\mathbf{I})^{-1})$. This problem arises in applications like Gaussian process regression.
@@
    We give faster algorithms for approximating Schatten $p$-norms and other matrix norms. For example, for the Schatten 1-norm (nuclear norm), we give an algorithm that runs in $\tilde{O}(n^{2.11})$ time, improving on an $\tilde{O}(n^{2.18})$ method of [Musco et al., ITCS 2018].
All results are proven in the real RAM model of computation. Interestingly, previous state-of-the-art algorithms for most of the problems above relied on stochastic iterative methods, like stochastic coordinate and gradient descent. Our work takes a completely different approach, instead leveraging tools from matrix sketching.
@@ 1 Introduction
We consider the complexity of solving a system of linear equations: given an $n\times n$ matrix $\mathbf{A}$ and an $n$-dimensional vector $\mathbf{b}$, find $\mathbf{x}$ such that $\mathbf{A}\mathbf{x}=\mathbf{b}$. This ubiquitous task in numerical linear algebra has applications across data science and machine learning, the physical sciences, engineering, and more. Despite many existing methods for solving linear systems, they remain a major computational bottleneck, so faster algorithms are a subject of active research. Much of the progress on faster algorithms concentrates on designing methods tailored for matrices with special structure, such as Laplacian, Toeplitz, or Hankel matrices, among others [KKM79, XXG12, ST14a, KMP12, KS16, CKK+18, PV21]. Progress on general, unstructured linear systems has been slower.
@@
When $k=n$, our fastest methods for solving 1.1 run in $O(n^{\omega})$ time via fast matrix multiplication. When $k=0$, the problem can be solved in $\tilde{O}(n^{2})$ time via any standard iterative solver such as CG. We are interested in the arguably more interesting intermediate regime, where $\mathbf{A}$ is a combination of a low-rank ill-conditioned matrix and a full-rank well-conditioned one. Such matrices arise often in practice, either due to various forms of regularization that are prevalent in machine learning, statistics and optimization [BV04, ZDW13, DW18, DM24], or due to the presence of isotropic noise coming from measurement error, rounding, or compression [LW11, LGW+21].
A standard approach to Problem 1.1 is to construct an approximation to the rank $k$ subspace identified by $\mathbf{A}$’s top singular values. That information can be used to precondition an iterative solver so that it runs in $\tilde{O}(n^{2})$ time [HMT11]. Constructing a sufficiently accurate approximation to the top subspace requires (block) power iteration or related methods [MM15], which take $\tilde{O}(n^{\omega(1,1,\log_{n}k)})$ time, where $\omega(1,1,\log_{n}k)\in[2,\omega]$ is the exponent of rectangular matrix multiplication between an $n\times n$ and $n\times k$ matrix [LG12]. [^1] This corresponds to the $O(n^{2}k)$ time for classical multiplication of $n\times n$ and $n\times k$ matrices. This approach already interpolates between the two extremes of the problem. In fact, it achieves a near-optimal $\tilde{O}(n^{2})$ runtime for $k=O(n^{0.32})$.
@@ 1.1 Main Results
$\tilde{O}\left(n^{2.065}\cdot\log^{3}1/\epsilon+k^{\omega}\right).$
@@
Surprisingly, our algorithm departs from the stochastic optimization approach used in the prior work of [DY24]. Instead, we employ a deterministic solver with a randomized preconditioner. To do so, we combine two main ingredients. First, building on prior work [FTU23, DEF+23], we show that even a very coarse approximation to $\mathbf{A}$’s top rank $k$ subspace can be used to construct a “good enough” preconditioner. In particular, we can replace the use of algorithms like block power method with more efficient, but less accurate, linear-time sketching methods for low-rank approximation [CW13]. Second, we show that it is possible to avoid the final step of such methods, which requires a prohibitive $\tilde{O}(nk^{\omega-1})$ cost to explicitly construct a rank $k$ subspace approximating $\mathbf{A}$’s top singular vectors. Instead, using $\tilde{O}(nk+k^{\omega})$ preprocessing time, we show how to maintain a data structure that lets us perform inexact matrix-vector products with an inverted low-rank preconditioner. The data structure involves additional levels of solving linear systems with randomized sketching and preconditioning methods, resulting in a framework that we call Multi-level Sketched Preconditioning (MSP). We discuss details of our approach further in Section 1.2. Before doing so, we highlight two additional applications of the framework beyond Theorem 1.1.
@@
An important case of linear systems that are well-conditioned except for a few large singular values are those that are explicitly regularized by adding a scaled identity, $\lambda\mathbf{I}$, to a positive semidefinite matrix $\mathbf{A}$. This setting arises, for instance, in kernel ridge regression [EAM14, RCR17, MM17, ACW17, FTU23] and second-order optimization [Lev44, Mar63, MS12, JJM23]. Prior results have shown that such linear systems can be solved in $\tilde{O}(n^{2}+n{d_{\lambda}}^{\omega-1})$ time, where $d_{\lambda}={\operatorname{tr}}(\mathbf{A}(\mathbf{A}+\lambda\mathbf{I})^{-1})\leq n$ is the so-called $\lambda$-effective dimension of the problem, often much smaller than $n$ [ACW17, MM17]. We improve on this by using Multi-level Sketched Preconditioning, leading to a time complexity of $\tilde{O}(n^{2}+{d_{\lambda}}^{\omega})$, which is optimal (up to log factors).
@@
$\tilde{O}\left(n^{2}\cdot\log^{3}1/\epsilon+{d_{\lambda}}^{\omega}\right),\quad\text{where}\quad d_{\lambda}={\operatorname{tr}}(\mathbf{A}(\mathbf{A}+\lambda\mathbf{I})^{-1}).$
@@
Linear systems often arise in algorithms for solving other matrix problems, including linear and semidefinite programming [CLS21, JKL+20], least squares and $l_{p}$ regression [Sar06, RT08, MM13, CP15, CD21], and for estimating various matrix properties [MNS+18]. Our new algorithms can be used to speed up any of these tasks. As a motivating example, we show how to improve methods for estimating various matrix norms, a fundamental problem in linear algebra. We build on an approach of [MNS+18], which uses regularized linear system solves to build rational function approximations that, when combined with stochastic trace estimation methods [Hut90], allow for the estimation of various functions of the singular values faster than $O(n^{\omega})$ time, i.e. faster than the time it takes to compute all singular values outright. [^3] Concretely, [MNS+18] estimates quantities of the form ${\operatorname{tr}}(f(\mathbf{A}))$, with the nuclear norm equal to the case where $f(x)=|x|$. Stochastic trace estimation methods like Hutchinson’s estimator and related techniques approximate ${\operatorname{tr}}(f(\mathbf{A}))$ by computing $\mathbf{g}^{T}f(\mathbf{A})\mathbf{g}$ for randomly chosen $\mathbf{g}$ [MMMW21].
@@ 1.2 Our Techniques
The starting point for our Multi-level Sketched Preconditioning comes from work on solving regularized systems of the form $(\mathbf{A}+\lambda\mathbf{I})\mathbf{x}=\mathbf{b}$, where $\mathbf{A}$ is positive definite (PD). Here, a central technique for preconditioning is the Nyström method [Nys30, WS01, GM16, MM17], which uses randomized sketching or sub-sampling to build a low-rank approximation for $\mathbf{A}$. As an illustration, let $\mathbf{S}\in\mathbb{R}^{l\times n}$ be a subsampling matrix, so that $\mathbf{A}\mathbf{S}^{\top}$ contains a random subset of $l$ columns from $\mathbf{A}$. The classical Nyström approximation is defined as $\hat{\mathbf{A}}_{\mathrm{nys}}=\mathbf{C}\mathbf{W}^{-1}\mathbf{C}^{\top}$, where $\mathbf{C}=\mathbf{A}\mathbf{S}^{\top}$ is the column submatrix and $\mathbf{W}=\mathbf{S}\mathbf{A}\mathbf{S}^{\top}$ is the corresponding $l\times l$ principal submatrix of $\mathbf{A}$. This can be extended to other sketching methods by replacing $\mathbf{S}$ with a sparse random matrix (CountSketch, OSNAP, LESS, etc. [CW13, NN13, Coh16, DLDM21]) or a Subsampled Randomized Hadamard Transform (SRHT, [AC09, Tro11]), all of which can be multiplied by $\mathbf{A}$ in $\tilde{O}(n^{2})$ time.
With sufficiently large sketch size (namely, proportional to the effective dimension $d_{\lambda}={\operatorname{tr}}(\mathbf{A}(\mathbf{A}+\lambda\mathbf{I})^{-1})$), one can show that $\mathbf{M}=\hat{\mathbf{A}}_{\mathrm{nys}}+\lambda\mathbf{I}$ is a good preconditioner for $\mathbf{A}+\lambda\mathbf{I}$, i.e., the condition number of $\mathbf{M}^{-1}(\mathbf{A}+\lambda\mathbf{I})$ is constant. At the same time, $\mathbf{M}^{-1}$ can be applied quickly due to its decomposition into $\mathbf{C}$ and $\mathbf{W}$. Concretely, using the inversion formula $\mathbf{M}^{-1}=\frac{1}{\lambda}\big{(}\mathbf{I}-\mathbf{C}(\mathbf{C}^{\top}\mathbf{C}+\lambda\mathbf{W})^{-1}\mathbf{C}^{\top}\big{)}$, we can compute $\mathbf{M}^{-1}\mathbf{r}$ for any vector $\mathbf{r}$ in $\tilde{O}(nl)$ time after precomputing $(\mathbf{C}^{\top}\mathbf{C}+\lambda\mathbf{W})^{-1}$ in $\tilde{O}(nl^{\omega-1})$ time.
@@
Average Condition Number Bounds via Nyström Preconditioning. We improve on standard Nyström preconditioning in two important ways. First, replacing the subsampling matrix $\mathbf{S}$ with a sparse sketching matrix [CNW16], we leverage a moment version of the oblivious subspace embedding property in conjunction with a spectral norm error low-rank approximation bound, to show that even when solving an unregularized system $\mathbf{A}\mathbf{x}=\mathbf{b}$, a preconditioner of the form $\mathbf{M}=\hat{\mathbf{A}}_{\mathrm{nys}}+\lambda\mathbf{I}$ can significantly improve conditioning. Note that the scalar $\lambda$ is now a parameter of the algorithm, and not of the problem, as was the case for standard Nyström preconditioning. We prove that, using Nyström approximation with sketch size $\tilde{O}(l)$ for any integer $l$, we can choose $\lambda$ so that the condition number of the preconditioned system $\mathbf{M}^{-1}\mathbf{A}$ can be reduced to $\tilde{O}(\frac{n}{l}\cdot\bar{\kappa}_{l})$, where $\bar{\kappa}_{l}=\frac{1}{n-l}\sum_{i>l}\frac{\sigma_{i}}{\sigma_{n}}$ is the average condition number of $\mathbf{A}$, excluding its top $l$ singular values (here, $\sigma_{i}$ are the singular values of $\mathbf{A}$ listed in a decreasing order). We observe that the quality of the preconditioner exhibits a dependence on $l$ both through the $\frac{n}{l}$ factor and the condition number $\bar{\kappa}_{l}$. Thus, obtaining the best condition number requires carefully tuning the parameter $\ell$ for each application. For example, if $\mathbf{A}$ has at most $k$ large singular values, as in 1.1, then $\bar{\kappa}_{l}=O(1)$ for any $l\geq k$, but due to the dependence on $\frac{n}{l}$, increasing $l$ past $k$ still improves the preconditioner. Unfortunately, this happens at the expense of increasing the $\tilde{O}(nl^{\omega-1})$ cost of inverting the Nyström approximation, which motivates our second contribution.
@@
Two-level Preconditioning for Positive Definite Systems. Instead of computing $\mathbf{M}^{-1}=\frac{1}{\lambda}\big{(}\mathbf{I}-\mathbf{C}(\mathbf{C}^{\top}\mathbf{C}+\lambda\mathbf{W})^{-1}\mathbf{C}^{\top}\big{)}$ explicitly, as in standard Nyström preconditioning, we focus on simply implementing matrix-vector multiplications with $\mathbf{M}^{-1}$, which suffice to apply a preconditioned iterative method like CG or Lanczos. The main challenge in doing so is to implement efficient multiplications with $(\mathbf{C}^{\top}\mathbf{C}+\lambda\mathbf{W})^{-1}$, i.e., to solve the system $(\mathbf{C}^{\top}\mathbf{C}+\lambda\mathbf{W})\mathbf{x}=\mathbf{r}$. We show how to do this using a second level of sketching and preconditioning. In particular, since $\mathbf{C}$ is a tall $n\times l$ matrix, the linear system $(\mathbf{C}^{\top}\mathbf{C}+\lambda\mathbf{W})\mathbf{x}=\mathbf{r}$ can be solved efficiently by preconditioning with $\mathbf{C}^{\top}{\mathbf{\Phi}}^{\top}{\mathbf{\Phi}}\mathbf{C}+\lambda\mathbf{W}$, where $\mathbf{\Phi}$ is an $O(l)\times n$ sketching matrix (specifically, an oblivious subspace embedding). The preconditioner $\mathbf{C}^{\top}{\mathbf{\Phi}}^{\top}{\mathbf{\Phi}}\mathbf{C}+\lambda\mathbf{W}$ can be constructed and inverted in just $\tilde{O}(nl+l^{\omega})$ time, which leads to an upfront cost of $\tilde{O}(nl+l^{\omega})$ for implementing $\tilde{O}(nl)$ time matrix-vector multiplications with $\mathbf{M}^{-1}$. This is a significant improvement on the $\tilde{O}(nl^{\omega-1})$ required by explicit inversion. Overall, we obtain a complexity of $\tilde{O}(n^{2}\sqrt{n/l}+l^{\omega})$ for solving 1.1 when $\mathbf{A}$ is positive definite. Optimizing over the parameter $l$ gives a final runtime bound of $\tilde{O}(n^{2.065}+k^{\omega})$ time for positive definite $\mathbf{A}$, which is our first step towards proving the general version of Theorem 3.2, which does not assume positive definiteness. See Section 4, Theorem 4.1, for details.
We note that to analyze the approach above, we leverage a stability analysis of the preconditioned Lanczos method, which ensures that an optimal convergence rate is still obtained even when the preconditioner $\mathbf{M}^{-1}$ is applied inexactly. We give this analysis in Section 6, building on prior work on the stability of the unpreconditioned Lanczos method [Pai71, Pai76, Gre89, MMS18]. Many alternative approaches would also suffice. For example, we could have obtained the same running times by using preconditioned Chebyshev iteration in place of Lanczos. The robustness of Chebyshev iteration to inexact applications of $\mathbf{M}^{-1}$ is well understood, and leveraged, e.g., in fast solvers for Laplacian systems [ST14b]. However, the method tends to converge much slower in practice than more popular methods like Lanczos or the Conjugate Gradient method. Thus, providing a stability analysis of the preconditioned Lanczos method is of independent interest beyond its application in our algorithms. Broadly, an increasing number of algorithms in numerical linear algebra combine iterative methods with inexact “inner loops”, often applied using randomized techniques. This approach has found applications in spectral density estimation [BKM22], quantum inspired linear algebra [BT24], and a number of other problems [OSV12]. All of this work hinges on stability analysis akin to our results on the preconditioned Lanczos method.
@@
Three-level Preconditioning for Indefinite Systems. The above strategy, which involves two levels of preconditioning, is restricted to positive definite matrices. Extending the approach to arbitrary linear systems requires additional work. A natural first attempt is to reduce an arbitrary $\mathbf{A}\mathbf{x}=\mathbf{b}$ linear system to a positive (semi-)definite linear system via the normal equations, $\mathbf{A}^{\top}\mathbf{A}\mathbf{x}=\mathbf{A}^{\top}\mathbf{b}$. However, if we apply the Nyström preconditioning approach discussed above, we realize that we require matrices $\mathbf{C}=\mathbf{A}^{\top}\mathbf{A}\mathbf{S}^{\top}$ and $\mathbf{W}=\mathbf{S}\mathbf{A}^{\top}\mathbf{A}\mathbf{S}^{\top}$, which can no longer be computed in $\tilde{O}(n^{2})$ time, even if $\mathbf{S}$ is a sparse random matrix. To address this, we show that the matrix $(\mathbf{C}^{\top}\mathbf{C}+\lambda\mathbf{W})^{-1}$ can nevertheless be applied efficiently by preconditioning with the approximation $(\mathbf{W}^{2}+\lambda\mathbf{W})^{-1}=\frac{1}{\lambda}\big{(}\mathbf{W}^{-1}-(\mathbf{W}+\lambda\mathbf{I})^{-1}\big{)}$. This preconditioner, in turn, can be applied by solving two linear systems associated with $\mathbf{W}^{-1}$ and $(\mathbf{W}+\lambda\mathbf{I})^{-1}$, both using a third level of sketching and preconditioning. Details are included in Section 5.